OpenStudy (anonymous):

When movies are being created, the set designers often use miniature models before they create the actual movie set. The set designers for Jurassic Park 3, created a 6 inch high and 4 inch long model dinosaur that represented a 20 feet high dinosaur on the set. How long was the actual dinosaur, rounded to the nearest foot? A. 15 feet B. 13 feet C. 14 feet D. 12 feet

OpenStudy (anonymous):

c

OpenStudy (anonymous):

how do you get that ?

OpenStudy (anonymous):

6/4=x/20. solve

OpenStudy (anonymous):

soo confused .. how do i solve that ?

OpenStudy (anonymous):

cross multiply and solve. 6 * 20= 4x

OpenStudy (anonymous):

120/4

OpenStudy (anonymous):

is 30

OpenStudy (anonymous):

my bad, wrong numbers

OpenStudy (anonymous):

but , that isn't in my options up there ): so now i really don't know what to do .

OpenStudy (anonymous):

6x=80

OpenStudy (anonymous):

what do i do for that ?

OpenStudy (anonymous):

so my answer would be 12 then right ? because i plugged in 2 for x and my answer is 12 . is that wrong ?

OpenStudy (anonymous):

80/6

OpenStudy (anonymous):

B is the answer.

OpenStudy (anonymous):

thank you very much ! (: